# Bondi mass, memory effect and balance law of polyhomogeneous spacetime

### v2

@@ 2. Asymptotic behaviour of Newman-Penrose quantities
$\begin{split}\bullet\ U=&U_{0}+\frac{U_{1}}{r}+\frac{U_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}U_{0}=0,\ ^{\#}U_{1}=N_{3}+1,\cdots,\\ &U_{0}=-\frac{1}{2},\\ &U_{1}=\sum_{j=0}^{N_{3}+1}U_{1,j}(\ln r)^{j},\\ &U_{1,N_{3}+1}=\gamma_{2,N_{3}+1}+\bar{\gamma}_{2,N_{3}+1},\\ &U_{1,j}=(j+1)U_{1,j+1}+\gamma_{2,j}+\bar{\gamma}_{2,j},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.38)
$\begin{split}\bullet\ X^{A}=&\frac{X_{2}^{A}}{r^{2}}+\cdots,\ \ \ \ ^{\#}X_{2}^{A}=N_{3},\cdots\\ &X_{2}^{A}=\sum_{j=0}^{N_{3}}X_{2,j}^{A}(\ln r)^{j},\\ &X_{2,N_{3}}^{A}=-\frac{1}{2}\bigg{[}\bar{\tau}_{2,N_{3}}\xi^{A}_{1,0}+\tau_{2,N_{3}}\bar{\xi}^{A}_{1,0}\bigg{]},\\ &X_{2,j}^{A}=\frac{1}{2}\bigg{[}(j+1)X^{A}_{2,j+1}-\bar{\tau}_{2,j}\xi^{A}_{1,0}-\tau_{2,j}\bar{\xi}^{A}_{1,0}\bigg{]},\ (j=N_{3}-1,\cdots,1,0).\end{split}$ (2.39)
$\begin{split}\bullet\ \omega=&\frac{\omega_{1}(z)}{r}+\frac{\omega_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\omega_{1}=N_{3}+1,\cdots,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ &\omega_{1}=\sum_{j=0}^{N_{3}+1}\omega_{1,j}(\ln r)^{j},\\ &\omega_{1,j+1}=-\frac{1}{j+1}(\bar{\alpha}_{2,j}+\beta_{2,j}),\ (j=0,1,\cdots,N_{3}),\\ &\omega_{1,0}=\bar{\eth}\sigma_{2,0}-\bar{\eth}\Psi_{0}^{3,0}+\bar{\eth}\sigma_{2,1}+2\omega_{1,2},\\ &\dot{\omega}_{1,0}=\bar{\eth}\dot{\sigma}_{2,0},\ \dot{\omega}_{1,2}=\frac{1}{2}\bar{\eth}\dot{\Psi}_{0}^{3,0}.\end{split}$ (2.40)
$\begin{split}\bullet\ \xi^{A}=&\frac{\xi^{A}_{1}}{r}+\frac{\xi_{2}^{A}(z)}{r^{2}}+\cdots,\ \ ^{\#}\xi^{A}_{1}=0,\ ^{\#}\xi^{A}_{2}=N_{3}+1,\cdots,\ \ \ \ \ \ \ \ \ \ \ \ \ \\ &\xi_{1}^{\theta}=\xi_{1,0}^{\theta}=\frac{1}{\sqrt{2}},\\ &\xi_{1}^{\varphi}=\xi_{1,0}^{\varphi}=\frac{i}{\sqrt{2}\sin\theta},\\ &\xi_{2}^{A}=\sum_{j=0}^{N_{3}+1}\xi_{2,j}^{A}(\ln r)^{j},\\ &\xi_{2,N_{3}+1}^{A}=-\sigma_{2,N_{3}+1}\bar{\xi}_{1,0}^{A},\\ &\xi_{2,j}^{A}=(j+1)\xi_{2,j+1}^{A}-\sigma_{2,j}\bar{\xi}_{1,0}^{A},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.41)
@@
$\begin{split}\bullet\ \Psi_{0}=&\frac{\Psi_{0}^{3}(z)}{r^{3}}+\frac{\Psi_{0}^{4}(z)}{r^{4}}+\cdots,\ \ \ \ ^{\#}\Psi_{0}^{3}=N_{3},\ ^{\#}\Psi_{0}^{4}=N_{4},\cdots,\\ &\frac{\partial\Psi_{0}^{3}}{\partial u}=0,\\ &\frac{\partial\Psi_{0}^{4}}{\partial u}=\frac{1}{2}(\Psi_{0}^{3})^{\prime}-\Psi_{0}^{3}+\eth\Psi_{1}^{3}.\end{split}$ (2.42)
$\begin{split}\bullet\ \Psi_{1}=&\frac{\Psi_{1}^{3}(z)}{r^{3}}+\frac{\Psi_{1}^{4}(z)}{r^{4}}+\cdots,\\ &{}^{\#}\Psi_{1}^{3}=N_{3},\ ^{\#}\Psi_{1}^{4}=\max\{N_{4}+1,2N_{3}+2\}\cdots,\\ &\Psi_{1}^{3}=\sum_{j=0}^{N_{3}}\Psi_{1}^{3,j}(\ln r)^{j},\\ &\Psi_{1}^{3,N_{3}}=\bar{\eth}\Psi_{0}^{3,N_{3}},\\ &\Psi_{1}^{3,j}=-(j+1)\Psi_{1}^{3,j+1}+\bar{\eth}\Psi_{0}^{3,j},\ (j=N_{3}-1,\cdots,1,0).\ \ \end{split}$ (2.43)
$\begin{split}\bullet\ \Psi_{2}=&\frac{\Psi_{2}^{3}}{r^{3}}+\cdots,\ \ \ \ ^{\#}\Psi_{2}^{3}=N_{3}+1,\dots\\ &\Psi_{2}^{3}=\sum_{j=0}^{N_{3}+1}\Psi_{2}^{3,j}(\ln r)^{j},\\ &\Psi_{2}^{3,j}=\frac{1}{j}\bigg{[}\bar{\eth}\Psi_{1}^{3,j-1}-\lambda_{1,0}\Psi_{0}^{3,j-1}\bigg{]},(j=1,2,\cdots,N_{3}+1),\\ &\dot{\Psi}_{2}^{3}=\eth\Psi_{3}^{2}+\sigma_{2}\Psi_{4}^{1},\\ &\frac{\partial\Psi_{2}^{3,0}}{\partial u}=-\eth^{2}\dot{\bar{\sigma}}_{2,0}-\sigma_{2,0}\ddot{\bar{\sigma}}_{2,0}.\end{split}$ (2.44)
$\begin{split}\bullet\ \Psi_{3}=&\frac{\Psi_{3}^{2}}{r^{2}}+\frac{\Psi_{3}^{3}}{r^{3}}+\cdots,\ \ \ ^{\#}\Psi_{3}^{2}=0,\ ^{\#}\Psi_{3}^{3}=N_{3}+1,\cdots,\\ &\Psi_{3}^{2}=\Psi_{3}^{2,0}=-\eth\dot{\bar{\sigma}}_{2,0},\\ &\Psi_{3}^{3}=\sum_{j=0}^{N_{3}+1}\Psi_{3}^{3,j}(\ln r)^{j},\\ &\Psi_{3}^{3,N_{3}+1}=-\bar{\eth}\Psi_{2}^{3,N_{3}+1},\\ &\Psi_{3}^{3,j}=(j+1)\Psi_{3}^{3,j+1}-\bar{\eth}\Psi_{2}^{3,j}+2\lambda_{1,0}\Psi_{1}^{3,j},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.45)
$\begin{split}\bullet\ \Psi_{4}=&\frac{\Psi_{4}^{1}}{r}+\frac{\Psi_{4}^{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\Psi_{4}^{1}=0,\ ^{\#}\Psi_{4}^{2}=0,\cdots,\\ &\Psi_{4}^{1}=-\ddot{\bar{\sigma}}_{2,0},\\ &\Psi_{4}^{2}=-\bar{\eth}\Psi_{3}^{2,0}=\bar{\eth}\eth\dot{\bar{\sigma}}_{2,0}.\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.46)
Here $\eth$ and $\bar{\eth}$ are the spin-weighted operators [30].
@@
3. Asymptotic symmetries of polyhomogeneous spacetimes
@@
$\begin{split}g_{00}&=1-\frac{2U_{1}}{r}+\cdots,\\ g_{01}&=1,\\ g_{02}&=\bigg{(}X^{\theta}_{2}-\frac{\omega_{1}+\bar{\omega}_{1}}{\sqrt{2}}\bigg{)}+\cdots,\\ g_{03}&=\bigg{(}X^{\varphi}_{2}+\frac{i(\omega_{1}-\bar{\omega}_{1})}{\sqrt{2}\sin\theta}\bigg{)}\sin^{2}\theta+\cdots,\\ \end{split}$ (3.4)
@@
$\begin{split}g_{11}&=0,\\ g_{12}&=0,\\ g_{13}&=0,\\ g_{22}&=-r^{2}+\sqrt{2}(\xi^{\theta}_{2}+\bar{\xi}^{\theta}_{2})r+\cdots,\\ g_{23}&=\frac{\sin\theta}{\sqrt{2}}\bigg{(}i\bar{\xi}^{\theta}_{2}-i\xi^{\theta}_{2}+\sin\theta(\xi^{\varphi}_{2}+\bar{\xi}^{\varphi}_{2})\bigg{)}r+\cdots,\\ g_{33}&=-r^{2}\sin^{2}\theta-i\sqrt{2}\sin^{3}\theta(\xi^{\varphi}_{2}-\bar{\xi}^{\varphi}_{2})r+\cdots.\\ \end{split}$ (3.5)
@@
where $f(\theta,\varphi)$ is an arbitrary smooth function on the 2-sphere, and $Y^{A}$ depends only on $\theta$ and $\varphi$ satisfying
$\mathcal{L}_{Y}q_{AB}=\frac{1}{2}(D_{C}Y^{C})\gamma_{AB}$ (3.8)
with $q_{AB}$ the round metric on the unit 2-sphere. Equation (3.8) implies that $Y^{A}$ is actually a conformal Killing vector field on the round sphere. Moreover, if $\xi_{1}^{a}$ and $\xi^{a}_{2}$ are both asymptotic infinitesimal symmetries, i.e., the components of $\xi_{1}^{a}$ and $\xi^{a}_{2}$ admit the form (3.7). Direct calculation shows that $[\xi_{1},\xi_{2}]^{a}$ is also an asymptotic infinitesimal symmetry. The above results are exactly the same as those obtained in the smooth asymptotically flat spacetime.
4. Bondi mass of smooth asympotically spacetime
@@
For the theory of general relativity, the Lagrange 4-form reads [19]
@@
Iyer and Wald defined the canonical energy as [19, Eq.(83)]
@@
Within the Bondi-Sachs coordinate system, the line element of an asymptotically flat spacetime is given by [29, Eq.(2.1)]
@@
Here $D$ stands for the covariant derivative of the standard unit 2-sphere. Therefore,
@@
Furthermore, the evolution equations of $m$ and $N_{A}$ are given by [29, Eqs.(3.1) and (3.4)]
@@
$\bullet\ \delta g_{uu}\sim O(\frac{1}{r}),$ (4.11)
$\bullet\ \delta g_{ru}\sim O(\frac{1}{r^{2}}),$ (4.12)
$\bullet\ \delta g_{rr},\delta g_{r\theta},\delta g_{r\varphi}=0,$ (4.13)
$\bullet\ \delta g_{u\theta},\delta g_{u\varphi}\sim O(1),$ (4.14)
$\bullet\ \delta g_{\theta\theta},\delta g_{\theta\varphi},\delta g_{\varphi\varphi}\sim O(r).$ (4.15)
@@
$=$ $-\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}g^{ce}\nabla_{e}(\frac{\partial}{\partial u})^{d}$
$=$ $-\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}g^{ce}\Gamma^{d}_{\ ef}(\frac{\partial}{\partial u})^{f}$
$=$ $-\frac{1}{16\pi}\int_{\infty}\epsilon_{abcd}g^{ce}\Gamma^{d}_{\ e0}$
$=$ $-\frac{1}{16\pi}\int_{\infty}\epsilon_{cdab}g^{ce}\Gamma^{d}_{\ e0}$
$=$ $-\frac{1}{16\pi}\int_{\infty}\big{(}\epsilon_{01ab}g^{0e}\Gamma^{1}_{\ e0}+\epsilon_{10ab}g^{1e}\Gamma^{0}_{\ e0}\big{)}$
@@
$\begin{split}\int_{\infty}\xi^{a}\Theta_{abc}&=\frac{1}{16\pi}\int_{\infty}(\frac{\partial}{\partial u})^{a}\epsilon_{dabc}g^{de}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int_{\infty}\epsilon_{0dbc}g^{de}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\ g^{re}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\ g^{re}g^{fh}(\partial_{f}\delta g_{eh}-\Gamma^{a}_{fe}\delta g_{ah}-\Gamma^{a}_{fh}\delta g_{ea}\\ &\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\partial_{e}\delta g_{fh}+\Gamma^{a}_{ef}\delta g_{ah}+\Gamma^{a}_{eh}\delta g_{fa})\\ &=\frac{-1}{16\pi}\int_{\infty}\mathrm{d}S\ g^{1e}g^{fh}(\partial_{f}\delta g_{eh}-\Gamma^{a}_{fh}\delta g_{ea}-\partial_{e}\delta g_{fh}+\Gamma^{a}_{eh}\delta g_{fa}).\end{split}$ (4.22)
Let
$(\rm{I})$ ::=:= $g^{1e}g^{fh}\partial_{f}\delta g_{eh},$ (4.23)
$(\rm{II})$ ::=:= $g^{1e}g^{fh}\Gamma^{a}_{\ fh}\delta g_{ea},$ (4.24)
$(\rm{III})$ ::=:= $g^{1e}g^{fh}\partial_{e}\delta g_{fh},$ (4.25)
$(\rm{IV})$ ::=:= $g^{1e}g^{fh}\Gamma^{a}_{\ eh}\delta g_{fa}.$ (4.26)
Then
$\int_{\infty}\xi^{a}\Theta_{abc}=-\frac{1}{16\pi}\int\bigg{[}({\rm{I})-(\rm{II})-(\rm{III})+(\rm{IV}})\bigg{]}\mathrm{d}S.$ (4.27)
@@
and it follows that
$g^{00}=g^{0A}=0,$ (4.33)
$g^{01}=-1-\frac{1}{16r^{2}}C_{AB}C^{AB}+\cdots,$ (4.34)
$g^{11}=1-\frac{2m}{r}+\cdots,$ (4.35)
$g^{1A}=\frac{q^{AB}D^{E}C_{BE}}{r^{2}}+\cdots,$ (4.36)
$g^{22}=\frac{1}{r^{2}}-\frac{C_{\theta\theta}}{r^{3}}+\cdots,$ (4.37)
$g^{23}=-\frac{C_{\theta\varphi}}{r^{3}\sin^{2}\theta}+\cdots,$ (4.38)
$g^{33}=\frac{1}{r^{2}\sin^{2}\theta}+\frac{C_{\theta\theta}}{r^{3}\sin^{2}\theta}+\cdots.$ (4.39)
In the following, we proceed to compute the terms $(\rm{I}),(\rm{II}),(\rm{III})$ and $(\rm{IV})$ respectively. For the term $(\rm{I})$, we have
$\begin{split}(\rm{I})&=g^{1e}g^{fh}\partial_{f}\delta g_{eh}\\ &=g^{10}g^{fh}\partial_{f}\delta g_{0h}+g^{11}g^{fh}\partial_{f}\delta g_{1h}+g^{1A}g^{fh}\partial_{f}\delta g_{Ah}\\ &=(\rm{I.1})+(\rm{I.2})+(\rm{I.3}),\end{split}$ (4.40)
where
$\begin{split}(\rm{I.1})&=g^{10}g^{fh}\partial_{f}\delta g_{0h}\\ &=g^{10}\big{[}g^{f0}\partial_{f}\delta g_{00}+g^{f1}\partial_{f}\delta_{01}+g^{fB}\partial_{f}\delta g_{0B}\big{]}\\ &=g^{10}\big{[}(g^{10}\partial_{1}\delta g_{00})+(g^{01}\partial_{0}\delta g_{01}+g^{11}\partial_{1}\delta g_{01}+g^{A1}\partial_{A}\delta g_{01})\\ &\ \ \ +(g^{1B}\partial_{1}\delta g_{0B}+g^{AB}\partial_{A}\delta g_{0B})\big{]}\\ &=\partial_{r}\delta g_{00}+\partial_{u}\delta g_{01}-g^{AB}\partial_{A}\delta g_{0B}+o(\frac{1}{r^{2}})\\ &=\partial_{r}\delta g_{00}+\partial_{u}\delta g_{01}-\frac{1}{r^{2}}\partial_{\theta}\delta g_{0\theta}-\frac{1}{r^{2}\sin^{2}\theta}\partial_{\varphi}\delta g_{0\varphi}+o(\frac{1}{r^{2}}),\end{split}$ (4.41)
${(\rm{I.2})}=g^{11}g^{fh}\partial_{f}\delta g_{1h}=g^{11}g^{f0}\partial_{f}\delta g_{10}=g^{11}g^{10}\partial_{r}\delta g_{10}=o(\frac{1}{r^{2}}),$ (4.42)
and
$\begin{split}{(\rm{I.3})}&=g^{1A}g^{fh}\partial_{f}\delta g_{Ah}\\ &=g^{1A}[g^{f0}\partial_{f}\delta g_{A0}+g^{fB}\partial_{f}\delta g_{AB}]\\ &=g^{1A}[g^{10}\partial_{1}\delta g_{A0}+g^{1B}\partial_{1}\delta g_{AB}+g^{CB}\partial_{C}\delta g_{AB}]\\ &=o(\frac{1}{r^{2}}).\end{split}$ (4.43)
Therefore,
${(\rm{I})}=\partial_{r}\delta g_{uu}+\partial_{u}\delta g_{ur}-\frac{1}{r^{2}}\partial_{\theta}\delta g_{u\theta}-\frac{1}{r^{2}\sin^{2}\theta}\partial_{\varphi}\delta g_{u\varphi}+o(\frac{1}{r^{2}}).$ (4.44)
Similarly, it can be obtained that
@@
$\int_{\infty}\xi^{a}\Theta_{abc}$ $=$ $\delta\bigg{[}-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\bigg{(}-\frac{2m}{r^{2}}-\frac{1}{16r^{2}}\partial_{u}(C_{AB}C^{AB}))+\frac{4m}{r}\bigg{]}$ (4.53)

### v3

@@ 2 Asymptotic behaviour of Newman-Penrose quantities
$\begin{split}U=&U_{0}+\frac{U_{1}}{r}+\frac{U_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}U_{0}=0,\ ^{\#}U_{1}=N_{3}+1,\cdots,\\ &U_{0}=-\frac{1}{2},\\ &U_{1}=\sum_{j=0}^{N_{3}+1}U_{1,j}z^{j},\\ &U_{1,N_{3}+1}=\gamma_{2,N_{3}+1}+\bar{\gamma}_{2,N_{3}+1},\\ &U_{1,j}=(j+1)U_{1,j+1}+\gamma_{2,j}+\bar{\gamma}_{2,j},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.38)
$\begin{split}X^{A}=&\frac{X_{2}^{A}}{r^{2}}+\cdots,\ \ \ \ ^{\#}X_{2}^{A}=N_{3},\cdots\\ &X_{2}^{A}=\sum_{j=0}^{N_{3}}X_{2,j}^{A}z^{j},\\ &X_{2,N_{3}}^{A}=-\frac{1}{2}\bigg[\bar{\tau}_{2,N_{3}}\xi^{A}_{1,0}+\tau_{2,N_{3}}\bar{\xi}^{A}_{1,0}\bigg],\\ &X_{2,j}^{A}=\frac{1}{2}\bigg[(j+1)X^{A}_{2,j+1}-\bar{\tau}_{2,j}\xi^{A}_{1,0}-\tau_{2,j}\bar{\xi}^{A}_{1,0}\bigg],\ (j=N_{3}-1,\cdots,1,0).\end{split}$ (2.39)
$\begin{split}\omega=&\frac{\omega_{1}(z)}{r}+\frac{\omega_{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\omega_{1}=N_{3}+1,\cdots,\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \\ &\omega_{1}=\sum_{j=0}^{N_{3}+1}\omega_{1,j}z^{j},\\ &\omega_{1,j+1}=-\frac{1}{j+1}(\bar{\alpha}_{2,j}+\beta_{2,j}),\ (j=0,1,\cdots,N_{3}),\\ &\omega_{1,0}=\bar{\eth}\sigma_{2,0}-\bar{\eth}\Psi_{0}^{3,0}+\bar{\eth}\sigma_{2,1}+2\omega_{1,2},\\ &\dot{\omega}_{1,0}=\bar{\eth}\dot{\sigma}_{2,0},\ \dot{\omega}_{1,2}=\frac{1}{2}\bar{\eth}\dot{\Psi}_{0}^{3,0}.\end{split}$ (2.40)
$\begin{split}\xi^{A}=&\frac{\xi^{A}_{1}}{r}+\frac{\xi_{2}^{A}(z)}{r^{2}}+\cdots,\ \ ^{\#}\xi^{A}_{1}=0,\ ^{\#}\xi^{A}_{2}=N_{3}+1,\cdots,\ \ \ \ \ \ \ \ \ \ \ \ \ \\ &\xi_{1}^{\theta}=\xi_{1,0}^{\theta}=\frac{1}{\sqrt{2}},\\ &\xi_{1}^{\varphi}=\xi_{1,0}^{\varphi}=\frac{i}{\sqrt{2}\sin\theta},\\ &\xi_{2}^{A}=\sum_{j=0}^{N_{3}+1}\xi_{2,j}^{A}z^{j},\\ &\xi_{2,N_{3}+1}^{A}=-\sigma_{2,N_{3}+1}\bar{\xi}_{1,0}^{A},\\ &\xi_{2,j}^{A}=(j+1)\xi_{2,j+1}^{A}-\sigma_{2,j}\bar{\xi}_{1,0}^{A},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.41)
@@
$\begin{split}\Psi_{0}=&\frac{\Psi_{0}^{3}(z)}{r^{3}}+\frac{\Psi_{0}^{4}(z)}{r^{4}}+\cdots,\ \ \ \ ^{\#}\Psi_{0}^{3}=N_{3},\ ^{\#}\Psi_{0}^{4}=N_{4},\cdots,\\ &\frac{\partial\Psi_{0}^{3}}{\partial u}=0,\\ &\frac{\partial\Psi_{0}^{4}}{\partial u}=\frac{1}{2}(\Psi_{0}^{3})^{\prime}-\Psi_{0}^{3}+\eth\Psi_{1}^{3}.\end{split}$ (2.42)
$\begin{split}\Psi_{1}=&\frac{\Psi_{1}^{3}(z)}{r^{3}}+\frac{\Psi_{1}^{4}(z)}{r^{4}}+\cdots,\\ &{}^{\#}\Psi_{1}^{3}=N_{3},\ ^{\#}\Psi_{1}^{4}=\max\{N_{4}+1,2N_{3}+2\}\cdots,\\ &\Psi_{1}^{3}=\sum_{j=0}^{N_{3}}\Psi_{1}^{3,j}z^{j},\\ &\Psi_{1}^{3,N_{3}}=\bar{\eth}\Psi_{0}^{3,N_{3}},\\ &\Psi_{1}^{3,j}=-(j+1)\Psi_{1}^{3,j+1}+\bar{\eth}\Psi_{0}^{3,j},\ (j=N_{3}-1,\cdots,1,0).\ \ \end{split}$ (2.43)
$\begin{split}\Psi_{2}=&\frac{\Psi_{2}^{3}}{r^{3}}+\cdots,\ \ \ \ ^{\#}\Psi_{2}^{3}=N_{3}+1,\dots\\ &\Psi_{2}^{3}=\sum_{j=0}^{N_{3}+1}\Psi_{2}^{3,j}z^{j},\\ &\Psi_{2}^{3,j}=\frac{1}{j}\bigg[\bar{\eth}\Psi_{1}^{3,j-1}-\lambda_{1,0}\Psi_{0}^{3,j-1}\bigg],(j=1,2,\cdots,N_{3}+1),\\ &\dot{\Psi}_{2}^{3}=\eth\Psi_{3}^{2}+\sigma_{2}\Psi_{4}^{1},\\ &\frac{\partial\Psi_{2}^{3,0}}{\partial u}=-\eth^{2}\dot{\bar{\sigma}}_{2,0}-\sigma_{2,0}\ddot{\bar{\sigma}}_{2,0}.\end{split}$ (2.44)
$\begin{split}\Psi_{3}=&\frac{\Psi_{3}^{2}}{r^{2}}+\frac{\Psi_{3}^{3}}{r^{3}}+\cdots,\ \ \ ^{\#}\Psi_{3}^{2}=0,\ ^{\#}\Psi_{3}^{3}=N_{3}+1,\cdots,\\ &\Psi_{3}^{2}=\Psi_{3}^{2,0}=-\eth\dot{\bar{\sigma}}_{2,0},\\ &\Psi_{3}^{3}=\sum_{j=0}^{N_{3}+1}\Psi_{3}^{3,j}z^{j},\\ &\Psi_{3}^{3,N_{3}+1}=-\bar{\eth}\Psi_{2}^{3,N_{3}+1},\\ &\Psi_{3}^{3,j}=(j+1)\Psi_{3}^{3,j+1}-\bar{\eth}\Psi_{2}^{3,j}+2\lambda_{1,0}\Psi_{1}^{3,j},\ (j=N_{3},\cdots,1,0).\end{split}$ (2.45)
$\begin{split}\Psi_{4}=&\frac{\Psi_{4}^{1}}{r}+\frac{\Psi_{4}^{2}}{r^{2}}+\cdots,\ \ \ ^{\#}\Psi_{4}^{1}=0,\ ^{\#}\Psi_{4}^{2}=0,\cdots,\\ &\Psi_{4}^{1}=-\ddot{\bar{\sigma}}_{2,0},\\ &\Psi_{4}^{2}=-\bar{\eth}\Psi_{3}^{2,0}=\bar{\eth}\eth\dot{\bar{\sigma}}_{2,0}.\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \end{split}$ (2.46)
[…]
@@
3 Asymptotic symmetries of polyhomogeneous spacetimes
@@
[…]
@@
[…]
@@
where $f(\theta,\varphi)$ is an arbitrary smooth function on the 2-sphere, $D_{A}$ stands for the covariant derivative of the standard unit 2-sphere, and $Y^{A}$ depends only on $\theta$ and $\varphi$ satisfying
$\mathcal{L}_{Y}q_{AB}=\frac{1}{2}(D_{C}Y^{C})q_{AB}$ (3.8)
with $q_{AB}$ the round metric on the unit 2-sphere. Equation (3.8) implies that $Y^{A}$ is actually a conformal Killing vector field on the round sphere. Moreover, if $\xi_{1}^{a}$ and $\xi^{a}_{2}$ are both asymptotic infinitesimal symmetries, i.e., the components of $\xi_{1}^{a}$ and $\xi^{a}_{2}$ admit the form (3.7), direct calculation shows that $\xi_{1},\xi_{2}]^{a}$ is also an asymptotic infinitesimal symmetry. The above results are exactly the same as those obtained in the smooth asymptotically flat spacetime.
4 Bondi mass of smooth asympotically flat spacetimes
@@
For the theory of general relativity, the Lagrange 4-form reads Iyer94
@@
Iyer and Wald defined the canonical energy as (Iyer94, , Eq.(83))
@@
Within the Bondi-Sachs coordinate system $\{u,r,x^{A}\}(A=2,3)$, the line element of an asymptotically flat spacetime is given by (CWYcmp2021, , Eq.(2.1))
@@
Here, the indices are contracted, raised, and lowered with respect to the round metric on the unit sphere $q_{AB}$, $m=m(u,x_{A})$ is the mass aspect, $N_{A}=N_{A}(u,x_{A})$ is the angular aspect, and $C_{AB}=C_{AB}(u,x_{A})$ is the shear tensor. The news tensor $N_{AB}$ is defined as
$N_{AB}=\frac{\partial C_{AB}}{\partial u}.$ (4.9)
It follows from Eqs.(4.7) and (4.8) that
@@
Furthermore, the evolution equations of $m$ and $N_{A}$ are given by (CWYcmp2021, , Eqs.(3.1) and (3.4))
@@
$\delta g_{uu}\sim O(\frac{1}{r}),$ (4.12)
$\delta g_{ru}\sim O(\frac{1}{r^{2}}),$ (4.13)
$\delta g_{rr},\delta g_{r\theta},\delta g_{r\varphi}=0,$ (4.14)
$\delta g_{u\theta},\delta g_{u\varphi}\sim O(1),$ (4.15)
$\delta g_{\theta\theta},\delta g_{\theta\varphi},\delta g_{\varphi\varphi}\sim O(r).$ (4.16)
@@
$\begin{split}\int_{\infty}\xi^{a}\Theta_{abc}&=\frac{1}{16\pi}\int_{\infty}(\frac{\partial}{\partial u})^{a}\epsilon_{dabc}g^{de}g^{fh}(\nabla_{f}\delta g_{eh}-\nabla_{e}\delta g_{fh})\\ &=-\frac{1}{16\pi}\int\bigg[({\rm{I})-(\rm{II})-(\rm{III})+(\rm{IV}})\bigg]\mathrm{d}S,\end{split}$ (4.23)
where
$(\rm{I})$ $:=$ $g^{1e}g^{fh}\partial_{f}\delta g_{eh},$ (4.24)
$(\rm{II})$ $:=$ $g^{1e}g^{fh}\Gamma^{a}_{\ fh}\delta g_{ea},$ (4.25)
$(\rm{III})$ $:=$ $g^{1e}g^{fh}\partial_{e}\delta g_{fh},$ (4.26)
$(\rm{IV})$ $:=$ $g^{1e}g^{fh}\Gamma^{a}_{\ eh}\delta g_{fa}.$ (4.27)
@@
we proceed to compute the terms $(\rm{I}),(\rm{II}),(\rm{III})$ and $(\rm{IV})$ respectively.
Direct calculation shows
${(\rm{I})}=\partial_{r}\delta g_{uu}+\partial_{u}\delta g_{ur}-\frac{1}{r^{2}}\partial_{\theta}\delta g_{u\theta}-\frac{1}{r^{2}\sin^{2}\theta}\partial_{\varphi}\delta g_{u\varphi}+o(\frac{1}{r^{2}}),$ (4.33)
@@
$\int_{\infty}\xi^{a}\Theta_{abc}$ $=$ $\delta\bigg[-\frac{1}{16\pi}\int_{\infty}\mathrm{d}S\bigg(-\frac{2m}{r^{2}}-\frac{1}{16r^{2}}\partial_{u}(C_{AB}C^{AB}))+\frac{4m}{r^{2}}\bigg]$ (4.42)